OpenStudy (anonymous):

Choose the point-slope form of the equation below that represents the line that passes through the point (−1, 6) and has a slope of −3. A.y − 6 = −3x − 3 B.y − 6 = −3(x + 1) C.y = −3x + 3 D.3x + y = 3

OpenStudy (anonymous):

someone please help!

OpenStudy (lgbasallote):

point slope form is \[\huge y- y_1 = m(x-x_1)\] where: m = slope x1 = x coordinate of the given point y1 = y coordinate of the given point does that help?

OpenStudy (anonymous):

not really so much..im kind of confused :(

OpenStudy (lgbasallote):

okay...what's your slope?

OpenStudy (anonymous):

-3 ?

OpenStudy (lgbasallote):

yes. what is the x-coordinate of your point?

OpenStudy (anonymous):

umm...right there is where i get confused. :(

OpenStudy (anonymous):

is the x-cordinate (-1,6) ?

OpenStudy (lgbasallote):

 does that help?

OpenStudy (anonymous):

ooooooooo ok i got it :D then what do we do?

OpenStudy (lgbasallote):

now you have m = -3 x1 = -1 y1 = 6 agree?

OpenStudy (anonymous):

yes :D

OpenStudy (anonymous):

so its B?

OpenStudy (lgbasallote):

yes it's B. sorry for late reply

OpenStudy (anonymous):

its ok ":D and thank you sooo much!!